# Pruning for Protection: Increasing Jailbreak Resistance in Aligned LLMs Without Fine-Tuning

### v1

@@ 3.4 Benchmarking on Standard Tasks
It is known that aggressive pruning reduces an LLM’s overall abilities sun2023wanda and therefore, we need to benchmark the pruned models’ performances across various tasks to ensure that the pruned models are still capable. For this reason, we chose to benchmark the models on Huggingface’s Open LLM Leaderboard open-llm-leaderboard , which consists of 7 benchmarks (6 tasks and the average performance on these tasks). Additionally, we measured the pruned models’ perplexities on the WikiText dataset wikitext2016 and the effective context length using the AltQA dataset pal2023giraffe . The latter evaluates the models’ ability to retrieve numerical answers to questions given long documents, which in our case was roughly 2k tokens to accommodate the maximum context length of our models. High performance on this task would indicate that long-range dependencies are still captured and therefore that the increased length of jailbreaking prompts does not affect the understanding of the malicious task.
@@ 4 Results
Figure 3: The top row illustrates the comparison of pruning effects for each jailbreaking prompt, the middle row aggregates the results by categories, while the bottom row aggregates the results based on severity levels. Pruning 20% of the weights results in an increased refusal rate for both the Vicuna 1.3 and LLaMA-2 Chat models However, pruning 30% weights hurts the safety of LLaMA-2 and shows reduced improvement for Vicuna. In contrast, Mistral Instruct v0.2 shows small but inconsistent changes post-pruning.
@@ 4.1 Quantitative Evaluation
We quantitatively measured the models’ resistance to generating harmful content. We display the comparisons of the jailbreaking success rates for our various models in Figure 3, providing a more detailed breakdown of the results in the Supplementary Material. As a general trend, we note that across the 5 types of malicious tasks, the average jailbreaking success rate varies significantly for different models. Figure 4 shows a comparison of the three base models.
(a) Breakdown by Jailbreak.
(b) Breakdown by Category.
(c) Breakdown by Severity.
Figure 4: Comparison of jailbreaking resistance of LLaMA-2 Chat, Vicuna 1.3 and Mistral Instruct v0.2. LLaMA-2 Chat is the most resilient against jailbreaking attacks, while Mistral is the least resilient.
@@ 4.2 Qualitative Comparison
We also qualitatively looked into responses generated by all the models. Table 1 Shows one example response from the base model and the pruned models. We did not find a noticeable degradation in the quality of the responses generated by the pruned models. We noticed that all LLaMA-2 7B Chat model outputs, including the base model outputs, were less informative and less malicious for the most complex jailbreaking prompts, such as GAME and TOMNJERRY, and more informative and malicious for easier prompts, such as CHARACTER and KEVIN.
4.3 Full Pruning vs MLP Pruning
LLaMA-2 7B Chat model has 32 Transformer Decoder blocks (touvron2023LLaMA, ). Each of these decoder blocks contains linear layers for both the attention mechanism and for the feed-forward network. In this experiment, we explored the effects of two pruning strategies: pruning every linear layer versus selectively pruning the layers in the feed-forward networks, which are implemented as multilayer perceptrons (MLP). Evaluating the jailbreaking resistance for these different strategies revealed a notable difference, the results of which are displayed in Figure 5. The model pruned to a sparsity 20% across both attention and MLP layers demonstrated increased jailbreaking resistance compared to selectively pruning the MLP to the same sparsity.
(a) Breakdown by Jailbreak.
@@ 4.3 Full Pruning vs MLP Pruning
Figure 5: The effects of Full Pruning vs MLP only pruning for LLaMA-2 7B Chat. The fully pruned model is more resistant to jailbreaking prompts.
4.4 Benchmarking Evaluation
Figure 6 summarizes the results of our pruned LLaMA-2 7B Chat models across various benchmarks. Table 2 summarizes our findings for Llama-2 Chat model. The corresponding benchmark results for Vicuna 1.3 and Mistral Instruct v0.2 are provided in the supplementary materials. We observe that generally, the pruned models perform competitively with and sometimes even surpass the base model. Since we do not observe significant degradation in the pruned models’ capabilities for reasoning, context length, or language modeling, the increased jailbreaking resistance from pruning LLaMA-2 and Vicuna cannot be attributed to a decrease in understanding of the task.
Figure 6: Comparison of performance of the unpruned models and 10-30% pruned models across various benchmarks.
@@ 4.4 Benchmarking Evaluation
[…]
ARC 52.90 53.16 52.47 52.47
HellaSwag 78.55 78.26 77.91 76.58
MMLU 48.32 48.18 47.27 45.57
TruthfulQA 45.57 45.29 45.88 44.82
Winogrande 71.74 71.49 70.72 69.61
Average 59.42 59.28 58.85 57.81
[…]
5-Shot 7.35 18.42 17.82 17.06
[…]
AltQA 63.38 63.82 63.60 60.09
[…]
WikiTextwikitext2016 6.94 7.02 7.17 7.33
[…]
5 Attention Pattern Analysis at Inference Time
We analyze attention patterns to aid interpretability(rogers2021primer, ) of our safety results. Specifically, we look at attention scores from LLaMA-2-7B-Chat at inference time. First, we aim to analyze attention pattern trends across all evaluation prompts, as well as across different transformer layers and heads. To achieve this, we have devised several metrics to summarize the sharpness and distribution of attention patterns. Our focus is specifically on the differences between original task tokens and ‘jailbreak’ tokens. For every token (in every prompt, transformer layer, and attention head), we rank the other tokens to which it attends and define the following metrics:
  • IntraTaskRank: The sum of ranks corresponding to attention links between in-malicious-task tokens.
  • EndTaskRank: The sum of ranks corresponding to attention links between the last token in the sequence and in-malicious-task tokens.
Formally, consider a transformer model with $L$ layers and $H$ heads, as well as a jailbreak prompt $p$ with $N_{p}$ tokens. Now, let $T_{p}\subseteq\{1,2,3,\ldots,N_{p}\}$ be the set of indices such that their corresponding tokens originate from the original malicious task.
For each layer $l\in\{1,2,3,\ldots,L\}$, attention head $h\in\{1,2,3,\ldots,H\}$, prompt $p$, and token index $i\in\{1,2,3,\ldots,N\}$ we define the list of ranks $[r_{i,1}^{(l,h,p)},r_{i,2}^{(l,h,p)},\ldots r_{i,N}^{(l,h,p)}]$ derived by sorting the $i^{th}$ row in the attention matrices at layer $l$ and head index $h$ when processing prompt $p$. Then,
$\mathrm{IntraTaskRank}(l,h,p)$ $=\mathbb{E}_{l,h,p}\left[\sum_{i,j\in\{1,2,3,\ldots,N_{p}\}}r_{i,j}^{(l,h,p)}[i\in T\land j\in T]\right]$
$\mathrm{EndTaskRank}(l,h,p)$ $=\mathbb{E}_{l,h,p}\left[\sum_{i\in\{1,2,3,\ldots,N\}}r_{N,i}^{(l,h,p)}[i\in T]\right]$
Intuitively, the lower the value of either of these metrics, the more invariant transformer models are to the jailbreak method. Additionally, it is interesting to disentangle whether jailbreak invariance is more correlated with the way we process malicious task tokens, or more with which tokens receive the most attention when generating the beginning of a new sequence.
Figure 7: IntraTaskRank Difference between Pruned and Base Models.
(a) Breakdown by severity.
(b) Breakdown by category.
(c) Breakdown by jailbreak.
Figure 7: IntraTaskRank Difference between Pruned and Base Models.
Figure 8: EndTaskRank Difference between Pruned and Base Models.
(a) Breakdown by severity.
(b) Breakdown by category.
(c) Breakdown by jailbreak.
Figure 8: EndTaskRank Difference between Pruned and Base Models.
We present our results in Figure 7 and Figure 8 and break down how both metrics differ between base and pruned models by task severity, task category, and jailbreak method. Across task severity levels and categories, pruned models consistently process tokens from the original malicious task paying less attention to the jailbreak pretext than base models do. The higher the target sparsity percentage is, the higher the difference is. We find these metrics’ behavior depends heavily on the jailbreak method, and in particular on whether or not the original malicious task is appended to the jailbreak pretext or surrounded by it.
Finally, we focus on the example jailbreak prompt from Table 1. We compare the base and 20% pruned LLaMA-2-7B-Chat models, specifically examining the attention scores that illustrate how much the original task tokens attend to each other. For various thresholds, we calculate the number of scores that exceed each threshold and plot the differences in these counts between the pruned and base models. This procedure is conducted by aggregating over all attention heads across all layers, and is also repeated by specifically focusing on attention heads in either the first or the last layer.
Figure 9: Comparing the counts of attention scores above various thresholds in both the base and the 20% pruned LLaMA-2-7B-Chat models reveals that the pruned models pay more attention to other original task tokens while processing tokens from the original malicious task prompt. The effect becomes significantly more pronounced in the later layers of the transformer.
(a) All layers.
(b) First layer.
(c) Last layer.
Figure 9: Comparing the counts of attention scores above various thresholds in both the base and the 20% pruned LLaMA-2-7B-Chat models reveals that the pruned models pay more attention to other original task tokens while processing tokens from the original malicious task prompt. The effect becomes significantly more pronounced in the later layers of the transformer.
We present our results in Figure 9 and confirm that the 20% pruned model pays less attention to jailbreak information in this successful case study.
Overall, we conclude that sharper attention patterns help the model resist jailbreak pretext scenarios. Pruning linear layers consistently provides safety improvements across a variety of tasks and severity levels, and they can be explained by numerous metrics that capture model in-variance to jailbreak token information. Albeit somewhat brittle to original malicious task token location in jailbreak, our method consistent safety improvements accross a variety of jailbreak methods used in practice. Additionally, note that this brittleness does not correlate with any decreased performance due to effects on effective context length, as underscored by our experiments on AltQA.
6 Conclusion
In this work, we explored the effects of pruning on the jailbreaking resistance of large language models. By applying Wanda pruning at varying levels of sparsity to LLaMA-2-7B-Chat, Vicuna 1.3 and Mistral Instruct v0.2 models, we obtained an assortment of compressed models with which to compare the susceptibilities of jailbreaking. We further curated a dataset of 225 malicious tasks and 2250 jailbreaking prompts for a total of 2475 prompts, with which we evaluated our base and compressed models. Our results show that if the unpruned model is sufficiently safety trained, then safety improves at lower sparsities of pruning, but then a reversal in the trend when pruned more aggressively. This suggests the possibility of using a carefully selected amount of pruning to aid in the deployment of safe LLMs.
For future directions to take with this work, we suggest a more comprehensive analysis of both base models and compression techniques. We primarily investigated Wanda pruning of 7-billion parameter models. However, it would be prudent to check whether these trends hold for larger models. Similarly, we chose this compression technique for its high efficacy and ease of usage, but exploring other means of compressing would provide a more robust understanding of the effects on safety. Overall, we have demonstrated that compression can have complicated impacts on the alignment of LLMs, and with the increasing push for scalable deployment of these models, it becomes all the more important that these consequences are well understood.
Acknowledgments and Disclosure of Funding
We thank MIT HAN Lab for their support in this work. ChatGPTchatgpt was used during the research and in preparation of this manuscript.
Supplementary Material
Here, we provide a detailed report of our evaluation results, broken down by jailbreaking category, model, and malicious task type.
Table 3: Refusal Rate for the LLaMA-2-7B-Chat models. All rates are displayed as percentages.
Jailbreak Category Model Hate Misinfo. Security Substance Unlawful
Unpruned 100.00 55.56 93.33 95.56 84.44
10% Pruned 100.00 53.33 91.11 95.56 86.67
ORIGINAL 20% Pruned 100.00 66.67 93.33 95.56 91.11
30% Pruned 100.00 55.56 93.33 95.56 88.89
Unpruned 64.44 44.44 73.33 75.56 51.11
10% Pruned 75.56 60.00 84.44 86.67 62.22
AIM 20% Pruned 82.22 68.89 88.89 95.56 82.22
30% Pruned 60.00 37.78 66.67 91.11 55.56
Unpruned 62.22 31.11 62.22 75.56 60.00
10% Pruned 64.44 31.11 73.33 82.22 57.78
CHARACTER 20% Pruned 66.67 35.56 75.56 82.22 57.78
30% Pruned 55.56 26.67 57.78 77.78 53.33
Unpruned 71.11 22.22 68.89 66.67 46.67
10% Pruned 68.89 31.11 73.33 64.44 53.33
CODE 20% Pruned 68.89 35.56 71.11 71.11 51.11
30% Pruned 46.67 15.56 33.33 55.56 31.11
Unpruned 53.33 6.67 51.11 57.78 40.00
10% Pruned 57.78 17.78 66.67 64.44 55.56
DAN 20% Pruned 57.78 17.78 64.44 71.11 48.89
30% Pruned 57.78 20.00 64.44 73.33 55.56
Unpruned 46.67 2.22 42.22 40.00 22.22
10% Pruned 51.11 2.22 37.78 53.33 20.00
GAME 20% Pruned 42.22 0.00 33.33 44.44 13.33
30% Pruned 22.22 0.00 6.67 17.78 4.44
Unpruned 68.89 40.00 84.44 80.00 77.78
10% Pruned 66.67 35.56 84.44 80.00 84.44
KEVIN 20% Pruned 71.11 31.11 91.11 88.89 82.22
30% Pruned 71.11 35.56 86.67 88.89 68.89
Unpruned 73.33 40.00 84.44 93.33 71.11
10% Pruned 86.67 31.11 82.22 86.67 73.33
PERSONGPT 20% Pruned 95.56 62.22 95.56 88.89 77.78
30% Pruned 93.33 62.22 93.33 91.11 91.11
Unpruned 93.33 75.56 91.11 97.78 77.78
10% Pruned 100.00 82.22 93.33 100.00 86.67
SIM 20% Pruned 100.00 95.56 97.78 100.00 95.56
30% Pruned 97.78 60.00 88.89 93.33 91.11
Unpruned 62.22 17.78 62.22 64.44 33.33
10% Pruned 62.22 17.78 60.00 57.78 33.33
TOMNJERRY 20% Pruned 57.78 17.78 64.44 55.56 48.89
30% Pruned 28.89 6.67 24.44 13.33 11.11
Unpruned 100.00 100.00 100.00 97.78 100.00
10% Pruned 100.00 100.00 100.00 100.00 100.00
TRANSLATOR 20% Pruned 100.00 100.00 100.00 100.00 100.00
30% Pruned 100.00 100.00 100.00 100.00 100.00
We also provide tables of benchmark results for Mistral Instruct v0.2 and Vicuna 1.3.
@@ 6 Conclusion
[…]
ARC 63.14 62.88 62.29 63.31
HellaSwag 84.88 84.85 84.90 84.37
MMLU 60.78 60.87 60.63 60.24
TruthfulQA 68.26 67.93 67.66 66.28
Winogrande 77.19 77.51 77.66 78.06
GSM8K 40.03 38.82 40.11 39.42
Average 65.71 65.48 65.54 65.28
[…]
AltQA 64.47 65.35 67.98 64.70
[…]
WikiText[16] 5.938 5.941 5.963 6.041
[…]
Pruned Sparsity
Benchmark Base 10% 20% 30%
[…]
ARC 50.43 51.45 52.67 51.11
HellaSwag 76.92 76.98 76.71 75.72
MMLU 48.14 47.95 47.27 46.54
TruthfulQA 47.01 46.88 47.22 45.53
Winogrande 70.48 69.77 69.06 68.98
Average 58.60 58.61 58.59 57.58
[…]
5-Shot 5.69 13.12 11.30 10.54
[…]
WikiText[16] 6.902 6.958 7.142 7.288
[…]

### v2

@@ 3.4 Benchmarking on Standard Tasks
It is known that aggressive pruning reduces an LLM’s overall abilities (Sun et al., 2023) and therefore, we need to benchmark the pruned models’ performances across various tasks to ensure that the pruned models are still capable. For this reason, we chose to benchmark the models on Huggingface’s Open LLM Leaderboard (Beeching et al., 2023), which consists of 6 tasks. Appendix C contains the description of the benchmarks. Additionally, we measured the pruned models’ perplexities on the WikiText dataset (Merity et al., 2016) and the effective context length using the AltQA dataset (Pal et al., 2023). The latter evaluates the models’ ability to retrieve numerical answers to questions given Wikipedia documents truncated to roughly 2k tokens each. The numerical answer for each document is modified to a different number to prevent the model from answering with pre-trained knowledge. High performance on this task would indicate that the effective context length is still intact after pruning. All the pruned models we tested performed nearly as well as the unpruned models. Since all jailbreaking prompts in our dataset are far smaller than 2k tokens, the safety enhancement in the pruned models cannot be explained by a degradation in effective context length.
@@ 4.1 Quantitative Evaluation
Table 1: Performance of different compressed models on key benchmarks from the Open LLM Leaderboard(Beeching et al., 2023) and on the AltQA(Pal et al., 2023) 2k-token benchmark. Scores excluding perplexity are presented in %. The base model is dense FP16 LLaMA-2-7B-Chat. For all benchmarks except perplexity, a higher score is better.
Pruned Sparsity
Benchmark Base 10% 20% 30%
ARC (25-shot) 52.90 52.90 53.41 53.41
HellaSwag (5-shot) 78.55 78.18 77.91 76.87
MMLU (5-shot) 48.32 48.10 47.49 47.04
TruthfulQA (6-shot) 45.57 45.40 45.84 45.02
Winogrande (5-shot) 71.74 71.43 70.72 71.03
GSM8K (0-Shot) 19.71 17.82 18.20 15.47
AltQA (0-shot) 52.19 52.63 51.97 48.68
Perplexity
WikiText(Merity et al., 2016) 6.943 7.019 7.158 7.259
We quantitatively measured the models’ resistance to generating harmful content. We display the comparisons of the jailbreaking success rates for our various models in Figure 3, Figure 7, and Figure 8. As a general trend, we note that across the 5 types of malicious tasks, the average jailbreaking success rate varies significantly for different models.
@@ 4.2 Qualitative Comparison
We also qualitatively looked into responses generated by all the models. Figure 2 Shows one example response from the base model and the pruned models. We did not find a noticeable degradation in the quality of the responses generated by the pruned models. We noticed that all model outputs, including the base model outputs, were less informative and less malicious for the most complex jailbreaking prompts, such as GAME and TOMNJERRY, and more informative and malicious for easier prompts, such as CHARACTER and KEVIN.
4.3 Benchmarking Evaluation
Table 1 summarizes our findings for Llama-2 Chat model. The corresponding benchmark results for Vicuna 1.3 and Mistral Instruct v0.2 are provided in Appendix C. We observe that generally, the pruned models perform competitively with and sometimes even surpass the base model. Since we do not observe significant degradation in the pruned models’ capabilities for reasoning, context length, or language modeling, the increased jailbreaking resistance from pruning LLaMA-2 and Vicuna cannot be attributed to a decrease in understanding of the task.
5 Interpretability
5.1 Pruning sharpens attention patterns
We inspect attention patterns and qualitatively observe that pruned models have sharper attention. Vig & Belinkov (2019) found that the entropy of attention patterns correlates with high-level semantic behavior: across various model depths, both the entropy of the attention patterns and their role in understanding sequence semantics evolve. Following this work, we calculate the entropy of attention patterns and average it over all prompts in our harmful tasks dataset, across layers and attention heads. In Figure 4, we illustrate the difference in average entropies between base and pruned models, noting that this reduction in average entropy reaches a plateau at a 20% prune percentage.
Figure 4: Difference of attention pattern entropies between base and pruned models. The pruned models demonstrate sharper attention patterns.
5.2 Sharper attention focuses on malicious tokens
Building on the observation that pruned models exhibit sharper attention patterns, we further analyze the distribution of attention across tokens. We measure the extent to which non-malicious ‘jailbreak’ tokens distract the model from focusing on malicious tokens. Following Vig & Belinkov (2019), we introduce a metric to capture the proportion of total attention that malicious tokens direct towards fellow malicious tokens. For every tokenized prompt $x$ in our dataset $\mathcal{X}$, we perform one forward pass and collect attention patterns $\alpha^{(l,h)}$ for every layer $l$ and attention head $h$. For a tokenized prompt $x$, we denote the set of indices originating from the original malicious task $\mathcal{T}_{x}$, while the remaining indices correspond to the different jailbreak pretexts. We introduce:
[…]
$\frac{\sum_{x\in\mathcal{X}}\sum_{l,h}\sum_{i=1}^{|x|}\sum_{j=1}^{i}\alpha_{ij}^{(l,h)}\llbracket i\in\mathcal{T}_{x},j\in\mathcal{T}_{x}\rrbracket}{\sum_{x\in\mathcal{X}}\sum_{l,h}\sum_{i=1}^{|x|}\sum_{j=1}^{i}\alpha_{ij}^{(l,h)}\llbracket i\in\mathcal{T}_{x}\rrbracket}$
This expression evaluates how effectively the model concentrates its attention on interactions among malicious tokens, despite the presence of distracting elements.
Figure 5: IgnoreJailbreak metric varies with the prune percentage, paralleling the safety refusal rate. This metric peaks at a pruning percentage of 20%, aligning with the peak of jailbreak resistance.
We present our results in Figure 5. We find that: i) pruning increases the IgnoreJailbreak metric; ii) IgnoreJailbreak peaks at a pruning percentage of 20%, corresponding with the peak in jailbreak resistance.
5.3 Perplexity Analysis
We now adopt an orthogonal approach to analyze, at a higher level of abstraction, how pruning influences language modeling capabilities. Our findings indicate that moderate pruning does not significantly impact language modeling performance on WikiText or the effective context length in AltQA tasks. However, this observation may not necessarily extrapolate to artificial constructs such as jailbreak templates. Indeed, it might even be preferable to have language models that do not overfit to such out-of-distribution prompts.
We approach this by investigating the perplexity assigned by both base and sparse models, to both the original malicious tasks and the prompts constructed using jailbreak templates. Note that model responses are not included in the following perplexity calculations.
In particular, for each original malicious task, we examine its perplexity before and after the application of jailbreak templates. For the latter, we report the perplexities associated with jailbreak attempts by calculating the average over the values obtained from the 10 jailbreak methods we examined.
Figure 6: Perplexity shifts when applying jailbreak templates to malicious prompts. Sparse models demonstrate a heightened capability to detect jailbreak templates compared to base models, assigning higher perplexity scores to original malicious tasks of equivalent perplexity levels.
We present our results in Figure 6. Sparse models consistently assign higher perplexity scores to jailbreak constructs than base models, when both assign similar perplexities to the corresponding original malicious tasks. This increased perplexity indicates that sparse models are more sensitive to deviations from the expected distribution of language, suggesting that WANDA acts as an effective regularizer. As demonstrated in Table 1, WANDA does not incur performance penalties when modeling in-distribution language passages. In contrast, it successfully detects out-of-distribution constructs.
6 Effects of WANDA Pruning on Linear Models with Correlated Input Features
In this section, we empirically validate that WANDA pruning significantly reduces test loss in Ordinary Least Squares (OLS) Regression models when the input features are correlated. This scenario is relevant in the context of large language models because natural language follows many structural patterns, such as power law, and the representations are not independent across different dimensions. Understanding the regularizing effects of WANDA pruning for a linear model can offer valuable insights for understanding its effects on more complex models.
Consider a set of inputs $X^{(d\times n)}$ with correlated features, true coefficients $w^{(1\times d)}$, and target $Y^{(1\times n)}$. Assume an i.i.d. white noise $\epsilon^{(1\times n)}\sim\mathcal{N}(0,\sigma^{2})$, leading to $Y=wX+\epsilon$. We take the ordinary least square (OLS) estimate of $w$ as $w^{\mathrm{OLS}}=((XX^{T})^{-1}XY^{T})^{T}$.
Let $X=(x^{(1)},\ldots,x^{(n)})$ and $Y=(y^{(1)},\ldots,y^{(n)})$, where $x^{(1)},\ldots,x^{(n)}$ are the input data points and $y^{(1)},\ldots,y^{(n)}$ are the corresponding outputs.
Define $w^{\mathrm{OLS}}=(w^{\mathrm{OLS}}_{1},\ldots,w^{\mathrm{OLS}}_{d})$. The WANDA pruning score for each $w^{\mathrm{OLS}}_{i}$ (where $d\geq i\geq 1$) is:
$s_{i}=|w^{\mathrm{OLS}}_{i}|\cdot\sqrt{\sum_{j=1}^{n}(x^{(j)}_{i})^{2}}$
In our experiments, we shall prune 30% of the weights of $w^{\mathrm{OLS}}$ with the smallest WANDA scores and observe the change in Mean Square Error (MSE) in test datasets.
We fix $w^{(1\times d)}$ and perform $N$ trials, each containing a training set $(X_{\mathrm{train}}^{(d\times n)},Y_{\mathrm{train}}^{(1\times n)})$ and a test set $(X_{\mathrm{test}}^{(d\times n)},Y_{\mathrm{test}}^{(1\times n)})$. All datasets share the same $w^{(1\times d)}$.
To generate a training dataset, first we sample a vector $\boldsymbol{\mathrm{x}}^{(1\times n)}\sim\mathcal{N}(0,1)$ and add perturbations $\delta^{(d\times n)}\sim\mathcal{N}(0,\alpha^{2})$ to it, resulting in $X_{\mathrm{train}}^{(d\times n)}=\boldsymbol{\mathrm{x}}^{(1\times n)}+\delta^{(d\times n)}$. The $\alpha$ controls the level of correlation in the input features. A low $\alpha$ indicates a high correlation among the input features and vice versa. After that, we sample $\epsilon^{(1\times n)}\sim\mathcal{N}(0,\sigma^{2})$ and create $Y_{\mathrm{train}}^{(1\times n)}=w^{(1\times d)}X_{\mathrm{train}}^{(d\times n)}+\epsilon^{(1\times n)}$. We sample another $\boldsymbol{\mathrm{x}}^{(1\times n)}$ and repeat the process for the test dataset.
Next, for each trial, we obtain $w^{\mathrm{OLS}}$ using the training samples, apply WANDA to prune 30% of the weights of $w^{\mathrm{OLS}}$, and then compare the MSE loss of the unpruned and the pruned estimators on the test dataset.
Our experiments involved $N=60$, $n=1000$, and we varied $d$ over $\{20,200,1000\}$, $\sigma$ over $\{0.2,0.6\}$, and $\alpha$ over $\{0.1,0.3\}$, resulting in a total of $3\times 2\times 2$ experimental settings. We performed a one-sample Z-test on the mean difference between the OLS estimator loss and the WANDA pruned estimator loss and reported the $p$-values. The WANDA pruned estimator consistently showed smaller MSE in the test dataset when the input features were highly correlated and irreducible error in the dataset was low. Table 2 summarizes our findings.
Table 2: Average test MSE loss comparison for $N=60$ trials. WANDA pruned estimator has a significantly smaller loss when the input features are highly correlated (small $\alpha$) and the irreducible error is low (small $\sigma$).
$d$ $\sigma$ $\alpha$ $\overline{\mathcal{L}_{\text{OLS}}}$ $\overline{\mathcal{L}_{\text{WANDA}}}$ $p$ value
20 0.2 0.1 1.48 1.45 $\ll 10^{-3}$
20 0.2 0.3 3.52 3.49 $\ll 10^{-3}$
20 0.6 0.1 2.50 2.56 $\sim 1.0$
20 0.6 0.3 24.30 24.24 0.004
200 0.2 0.1 262.35 262.27 $\ll 10^{-3}$
200 0.2 0.3 115.59 115.51 $\ll 10^{-3}$
200 0.6 0.1 92.68 92.64 0.012
200 0.6 0.3 27.55 27.53 $\ll 10^{-3}$
1000 0.2 0.1 364.36 363.25 0.004
1000 0.2 0.3 1298.23 1297.83 0.018
1000 0.6 0.1 119772.62 117906.12 0.088
1000 0.6 0.3 2004.06 1978.05 0.114
7 Conclusion
In this work, we explored the effects of pruning on the jailbreaking resistance of large language models. By applying Wanda pruning at varying levels of sparsity to LLaMA-2-7B-Chat, Vicuna 1.3, and Mistral Instruct v0.2 models, we obtained an assortment of compressed models with which to compare the susceptibilities of jailbreaking. We further curated a dataset of 225 malicious tasks and 2250 jailbreaking prompts, with which we evaluated our base and compressed models. Our results show that if the unpruned model is sufficiently safety trained, then safety improves at lower sparsities of pruning, but then a reversal in the trend when pruned more aggressively. This suggests the possibility of using a carefully selected amount of pruning to aid in the deployment of safe LLMs.
For future directions to take with this work, we suggest a more comprehensive analysis of both base models and compression techniques. We primarily investigated the WANDA pruning of 7-billion parameter models. However, it would be prudent to check whether these trends hold for larger models. Similarly, we chose this compression technique for its high efficacy and ease of usage, but exploring other means of compressing would provide a more robust understanding of the effects on safety.
8 Impact Statement
This paper presents work whose goal is to introduce a novel method for improving the resistance of LLMs to producing harmful content. The societal consequences of our work may, however, extend beyond our original intentions. Open-sourcing a dataset of jailbreaking prompts and malicious tasks could potentially enable the generation of more dangerous content. We acknowledge this risk, but with the rapid growth in usage of LLMs and the need for safety as well, we hope that our work and dataset will enable and promote further research in this area.
Appendix A Detailed Safety Results
Below we present the detailed safety results for Vicuna 1.3 and Mistral Instruct v0.2
(a) By Jailbreak.
@@ Appendix A Detailed Safety Results
Figure 7: Vicuna 1.3 7B shows moderate safety improvement post-pruning.
(a) By Jailbreak.
(b) By Category.
(c) By Severity.
Figure 8: Mistral Instruct v0.2 7B shows minimal safety improvement post-pruning.
Appendix B Attention Pruning vs Full Pruning vs MLP Pruning
In our study of the LLaMA-2 7B Chat model, which comprises 32 Transformer Decoder blocks (Touvron et al., 2023), we focused on three pruning strategies: pruning every attention layer, every linear layer and pruning the layers of the multi-layer perceptron (MLP). Evaluating the jailbreaking resistance for these different strategies revealed a notable difference, the results of which are displayed in Figure 9. Intriguingly, the model achieved the highest resistance to jailbreaking when pruned to 20% sparsity exclusively in the attention layers, outperforming both the selective MLP layer pruning and the uniform pruning across all layers.
(a) By Jailbreak.
(b) By Category.
(c) By Severity.
Figure 9: The effects of attention layer pruning vs full Pruning vs MLP-only pruning for LLaMA-2 7B Chat. The attention pruned model is the most resistant to jailbreaking prompts.
Appendix C Details about the Benchmarks
  • ARC (AI2 Reasoning Challenge): ARC is a benchmark consisting of grade-school level multiple-choice science questions, designed to assess a system’s ability to apply reasoning and understanding of basic scientific concepts. (Clark et al., 2018) It challenges AI models to go beyond pattern recognition and engage in elementary forms of reasoning.
  • HellaSwag: HellaSwag is a dataset aimed at testing the commonsense reasoning and contextual understanding of AI systems, where the task is to predict the correct ending to a given scenario among multiple choices, often requiring an understanding of implicit real-world knowledge. (Zellers et al., 2019)
  • MMLU: Massive Multitask Language Understanding (MMLU) is a comprehensive benchmark encompassing a wide range of subjects and domains, from humanities to natural sciences, intended to evaluate an AI model’s broad understanding and reasoning capabilities across diverse topics. (Hendrycks et al., 2021)
  • TruthfulQA: TruthfulQA is designed to assess the ability of language models to provide truthful and factual answers. (Lin et al., 2022) It consists of questions that are intentionally misleading or prone to the elicitation of falsehoods, testing the model’s resistance to propagating inaccuracies.
  • Winograde: The Winograde Schema Challenge is a natural language understanding test focusing on coreference resolution, where the task is to resolve ambiguity in sentences that require understanding the relationships between different entities. (Sakaguchi et al., 2019)
  • GSM8k: Grade School Math 8k (GSM8k) is a benchmark consisting of grade-school level math problems, designed to evaluate an AI’s capability in understanding and solving basic arithmetic and mathematical reasoning questions. (Cobbe et al., 2021)
  • AltQA: This benchmark evaluates the models’ ability to retrieve numerical answers to questions given Wikipedia documents truncated to roughly 2k tokens each. The numerical answer for each document is modified to a different number to prevent the model from answering with pre-trained knowledge. (Pal et al., 2023) High performance on this task would indicate that the effective context length is still intact after pruning.
  • Perplexity: Perplexity is a measurement used to assess the performance of language models, indicating how well a model predicts a sample; a lower perplexity score means the model is more confident and accurate in its predictions. Mathematically, it is defined as the exponentiated average negative log-likelihood of a sequence of words, given as $PP(W)=\sqrt[N]{\prod_{i=1}^{N}\frac{1}{P(w_{i}|w_{1},\ldots,w_{i-1})}}$, where $PP(W)$ is the perplexity of the word sequence $W$, $N$ is the length of the sequence, and $P(w_{i}|w_{1},\ldots,w_{i-1})$ is the probability of word $w_{i}$ given the preceding words.
Here we provide tables of benchmark results for Mistral Instruct v0.2 and Vicuna 1.3.
Table 3: Mistral Instruct v0.2 performance on 8 key benchmarks. Scores excluding perplexity are presented in %. For all benchmarks except perplexity, a higher score is better.
@@ Appendix C Details about the Benchmarks
ARC (25-shot) 63.14 63.05 62.88 62.97
HellaSwag (5-shot) 84.88 84.88 84.84 84.71
MMLU (5-shot) 60.78 60.84 60.81 60.49
TruthfulQA (6-shot) 68.26 68.11 68.26 67.49
Winogrande (5-shot) 77.19 77.11 77.90 77.98
GSM8K (0-Shot) 28.20 27.82 27.45 29.11
AltQA (0-shot) 58.77 58.99 60.31 57.46
Perplexity
WikiText(Merity et al., 2016) 5.938 5.938 5.941 5.944
Table 4: Vicuna 1.3 performance on 7 key benchmarks. Scores excluding perplexity are presented in %. Evaluation on the AltQA(Pal et al., 2023) 2k-token benchmark is omitted due to exceeding the maximum sequence length of the model.
@@
ARC (25-shot) 50.43 52.22 52.30 51.02
HellaSwag (5-shot) 76.92 77.05 77.05 76.41
MMLU (5-shot) 48.14 47.93 47.39 47.04
TruthfulQA (6-shot) 47.01 46.87 46.83 46.06
Winogrande (5-shot) 70.48 69.53 69.22 69.03
GSM8K (0-Shot) 6.37 5.99 5.91 4.55
AltQA (0-shot) - - - -
Perplexity
WikiText(Merity et al., 2016) 6.902 6.949 7.086 7.163
Appendix D ChatGPT System Prompt